OpenStudy (anonymous):

Part A: Create an example of a polynomial in standard form. How do you know it is in standard form? Part B: Explain the closure property as it relates to polynomials. Give an example. Will Medal And Fan

OpenStudy (anonymous):

Please Help

OpenStudy (anonymous):

@Mehek14

OpenStudy (anonymous):

ima fan real quick to you mehmek

Mehek (mehek14):

make any polynomial but in order from greatest exponent to the least

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

4xy^3+2x+5

Mehek (mehek14):

good so that's part A

Mehek (mehek14):

"How do you know it is in standard form? " put because the exponents are from greatest to least

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

how do i explain the closure property?

Mehek (mehek14):

whether you add/subtract/divide/multiply polynomials, the answer is always a polynomial

OpenStudy (anonymous):

and how do i give an example

Mehek (mehek14):

\((x+2)(x+1)\)

Mehek (mehek14):

\(x*x=x^2\\x*1=1x\\2*x=2x\\2*1=2\\x^2+3x+2\)

Mehek (mehek14):

the product is still a polynomial

OpenStudy (anonymous):

thank you

Mehek (mehek14):

yw :)